OpenStudy (anonymous):

What is the mean of the data set? 18 8 15 18 11

OpenStudy (anonymous):

Add them all together and divide by 5.

OpenStudy (y1066380):

Sum=18+8+15+18+11 Number of data=5 So, you will have (18+8+15+18+11)/5 =14

OpenStudy (arabpride):

yup, to find the mean: 18 + 8 + 15 + 18 + 11 / 5

OpenStudy (anonymous):

Ty

OpenStudy (arabpride):

soo.... 70/5 average = 14

OpenStudy (arabpride):

Welcome *^.^*